The divided line segment from A(-9,2) to point B(12,8) is divided by the point P(-2,4) in the ratio m:n. In what ratio does P partition AB?
Answer by Boreal(15235)

The distance formula is sqrt ((x1-x2)^2+(y1-y2)^2)
The AB distance is sqrt (21^2+6^2)=sqrt(441+36)=sqrt(477)
The AP distance is sqrt (49+4)=sqrt(53)
sqrt(53)/sqrt(477)=sqrt(53)/sqrt(9)(sqrt(53)
This is 1/sqrt(9)=1/3
it portions AB in a 1:2 ratio, where AP is (1/2) the distance to PB.
It is 1/3 of the way from A to B.